OpenStudy (anonymous):

x: -2, -1, 3, 4, 5, y: -7, -5, 3, 5, 7, find y when x=2

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

how yo u get that

OpenStudy (anonymous):

how you get that

OpenStudy (ybarrap):

-7 = -2*3 -1 = -2 - 5 -5 = -1*3 - 2 = -1 - 4 3 = 3*3 - 6 = 3 + 0 5 = 4*3 - 7 = 4 + 1 7 = 5*3 - 8 = 5 +2 missing add numbers in y -3 = 0 - 3 -1 = 1 - 2 1 = 2 - 1 pattern -(2n+1) = -(1-n) - (2n-1), where n = any integer If we let x = -(1-n) and y = -(2n+1), then y = x - (2n-1), for any given integer if x = 2, then x = -(1-n)=2 --> n = 3 and y = -(2n+1) --> y = -(2*3+1) = -7 so when x = 2, y = -7